OpenStudy (anonymous):

Use the quadratic formula to solve the equation, 2x^2-3x=-5...I came up with x=1/4(3-i√31

OpenStudy (anonymous):

\(\huge x=\frac{-b\pm\sqrt{b^2-4ac}}{2a} \) a=2 b=-3 c=5 so \(\huge x=\frac{-(-3)\pm\sqrt{(-3)^2-4(2)(5)}}{2(2)} \) \(\huge x=\frac{3\pm\sqrt{9-40}}{4} \) \(\huge x=\frac{3\pm \sqrt{-31}}{4} \) \(\huge x=\frac{3\pm i\sqrt{31}}{4} \)